### v2

@@ Abstract
Generating speech-driven 3D talking heads presents numerous challenges; among those is dealing with varying mesh topologies where no point-wise correspondence exists across all meshes the model can animate. While simplifying the problem, it limits applicability as unseen meshes must adhere to the training topology. This work presents a framework capable of animating 3D faces in arbitrary topologies, including real scanned data. Our approach relies on a model leveraging heat diffusion to predict features robust to the mesh topology. We explore two training settings: a registered one, in which meshes in a training sequences share a fixed topology but any mesh can be animated at test time, and an fully unregistered one, which allows effective training with varying mesh structures. Additionally, we highlight the limitations of current evaluation metrics and propose new metrics for better lip-syncing evaluation between speech and facial movements. Our extensive evaluation shows our approach performs favorably compared to fixed topology techniques, setting a new benchmark by offering a versatile and high-fidelity solution for 3D talking head generation where the topology constraint is dropped.
keywords:
3D Talking Heads, 3D Face Animation, Unregistered Mesh, Speech-Driven, Evaluation Metrics, Geometric Deep Learning.
@@ 1 Introduction
3D talking heads refer to the task of animating a generic 3D face mesh so that it corresponds to a speech, given in the form of an audio signal. This task has received significant attention in recent years, thanks to the wide range of potential applications, across diverse fields, that it can open. This includes computer-generated imagery (CGI) in movies, video games, virtual reality, and medical simulations. Nevertheless, it also faces unique challenges, primarily arising from the complexity of mapping the semantics of the spoken words contained in the audio to the intricacy of facial motions. A major challenge for 3D talking heads is that of generating convincing and audio-synchronized lip movements that faithfully replicate the spoken sentence. Several other aspects, though, contribute to making a generated animation realistic. For example, head movements, facial expressions or speech styles are all important factors, and numerous approaches were proposed to account for such diverse challenges [31, 36, 47]. In all cases, the primary objective is that of learning a mapping between the audio input to non-rigid and time-dependent deformations of the face mesh that allow for its animation.
@@
In this regard, all state-of-the-art methods in this field simplify the problem by assuming that the meshes to be animated are aligned to a fixed topology, that is, all face meshes share a consistent number and arrangement of points. On one hand, this constraint allows for eluding most difficulties related to processing unordered point-clouds. With the meshes in vertex-wise correspondence and their structure known and consistent, one can, for instance, use effective graph operators [4, 12, 19] or define specific losses to control localized face areas [50]. On the other hand, it puts several constraints on the applicability. For example, existing approaches require training separate models for each specific topology, hindering its generalization power. In addition, if one wants to animate a new mesh, either captured with a real scanner or designed with some graphics software, it is first required to process the mesh to make it consistent with the topology used for training. Finally, it also adds a level of complexity in the data collection since newly captured mesh sequences need to be processed in order to align them to a common topology, which unveils new challenges mostly when non-rigid face deformations are involved [16, 18].
@@
In addition to the aforementioned issues, another critical aspect is related to the evaluation protocols and metrics. Thoroughly assessing the performance of a 3D talking head method is complex, as it requires to simultaneously verify the extent to which: (i) the overall shape of the animated face is preserved. There is in fact no guarantee that mouth motions won’t negatively affect other face regions; (ii) the accuracy of the lip motion and synchronization is maintained, both in terms of shape similarity, perceived realism and motion dynamics. We identified a clear shortcoming in the current literature in this regard. In most works, the primary metrics used for quantitative evaluation and comparison measure the discrepancy between ground-truth and generated animation in terms of per-frame geometric discrepancy of vertices [10, 14, 35, 47]. These are referred to as Lip-Vertex Error (LVE), Mean Vertex Error (MVE) and Face Dynamics Deviation (FDD). We argue that these metrics do not provide a thorough picture of a method performance, and while these quantitative measures are widely used, they often fall short in providing a comprehensive evaluation. Suffice it to say, the evaluation of the motion dynamics and synchronization is largely ignored, and lower metrics do not necessarily imply a more realistic animation. A common workaround involves performing qualitative user studies, where people are asked to evaluate the quality of the generated samples. While useful to some extent, these studies lack a clear robust protocol, and usually involve a few users and examples, posing doubts on their reliability. Some papers raised similar concerns, and defined additional metrics to fill in such gap. For instance, in [35] the Lip Readability Percentage (LRP) is introduced. However, it is still based on per-frame lip vertex distances. Other examples are Dynamic Time Warping (DTW) used in Imitator [50] for assessing temporal coherence, or the Displacement Angle Discrepancy (DAE) proposed in [29], which provides a measure of the angular difference between subsequent frames. These measures, and others, are a clear evidence that the current benchmarking miss a piece, and that a stable, consistent and comprehensive set of complementary error measures needs to be defined for proper evaluation across methods.
@@
In summary, ScanTalk, illustrated in Figure 1, represents a significant advancement in the field of speech-driven 3D face animation by overcoming the limitations of fixed-topology models and establishing new standards for evaluation and training. Our framework demonstrates broad applicability across different datasets and conditions. Parts of the current work have been introduced in our previous paper [30]. The present extension provides the following new contributions:
  • In [30], ScanTalk can be trained on homogeneous sequences: each frame of the sequence shares the topology of the first frame. We modified the training scheme and defined appropriate losses so that the model can be trained with inhomogeneous sequences: each frame of the sequence has an arbitrary topology. By dropping this constraint, ScanTalk can be trained with truly unregistered meshes.
  • While the Chamfer distance is widely used as a loss for unsupervised static 3D reconstruction, we propose to use a dynamic extension of this loss for a complete unsupervised setting to learn speech-driven motion prediction.
@@ 2.1 Animation of Unregistered 3D Face Meshes
Animating a 3D face mesh to reproduce expressions or speech is a research topic that witnessed noticeable advancements only recently, thanks to the progresses achieved in the design of effective deep architectures for processing 3D data [4, 26, 32, 33]. However, it is well known that deep models are data-hungry, which poses the problem of collecting sufficiently large 4D datasets. Moreover, all state-of-the-art methods based on deep learning require the face meshes to be in a fixed topology.
In the realm of 3D face mesh acquisition for datasets collection, methodologies vary from extracting geometry from images or videos [58] or generating synthetic data [38], to employing specialized scanners in controlled environments [15, 17, 41, 56, 59]. While methods that follow the former paradigm are simpler and cost-effective, and typically result in meshes with known topology, they may not capture complete 3D information with the necessary fidelity. On the other hand, 3D scans, though rich in captured details of the face, often come unregistered. The varying mesh structure and nuisances resulting from the sensors, e.g., holes or noise, further complicate a direct animation. In response, a widely used approach to address this challenge is that of first registering the raw scans onto predefined topologies, typically by means of 3D Morphable Models [15, 16, 24, 28, 56].
While some deep learning approaches have emerged to handle face scans with unknown structure through latent representations with robust encoders [46, 54] built upon PointNet [5, 39] or adaptive convolution strategies [53], these methods still rely on a registered decoding strategy, which can smooth out important details. Recent advancements closer to our work include DiffusionNet [46] combined with neural Jacobian fields [1] in Neural Face Rigging (NFR) [40], which facilitates animation transfer between sequences of unregistered face meshes. Our framework stands out by introducing a model that leverages audio data to animate unregistered meshes directly, without fixed topological constraints.
@@
Numerous methodologies have emerged to synchronize facial animations with speech. While significant progress has been made in 2D talking heads [22, 52], extending these approaches to 3D faces remains challenging. Nevertheless, generating 3D talking heads from audio has seen significant advancements, initially pioneered by [10] and greatly improved through deep learning methods. Recent works leverage large datasets and deep learning strategies [10, 14, 20, 29, 31, 43, 47, 49, 50, 57]. VOCA [10] was among the first to develop a generalizable model for 3D talking heads but lacked expressiveness and convincing lip synchronization. MeshTalk [43] improved upon this with a larger registered dataset and a more expressive model. FaceFormer [14] utilized a transformer architecture and a pre-trained audio encoder, like Wav2vec2 [45], with further improvements in FaceXHubert [20] and FaceDiffuser [47], demonstrating enhanced cross-modality mappings from audio to facial motion. From a different perspective, Nocentini et al. [29] designed a coarse-to-fine approach named S2L-S2D, and showed that lip motion can be effectively modeled by learning to first animate facial landmarks, and then reconstructing the whole face shape. Recent models like CodeTalker [57] and SelfTalk [35] introduced new techniques to address over-smoothing and enhance lip motion expressiveness. Integrating emotions [11, 36] and head poses [48, 60] into models further pushes the boundaries, though the lack of registered data remains a significant challenge. In this regard, a recent work attempted to bridge this gap by proposing a data-driven approach to synthetically generate a large-scale dataset of expressive 3D talking heads, named EmoVOCA [31], which combines the VOCAset dataset [10] with 4D expressive sequences from the Florence4D dataset [38]. However, creating this dataset was made easy as both VOCAset and Florence4D contain meshes in the same topology.
@@
In 2D talking head generation, metrics like Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), are commonly employed to evaluate visual fidelity. Meanwhile, lip synchronization can be measured using metrics such as Lip Sync Error Confidence (LSE-C), and Lip Sync Error Distance (LSE-D), which rely on visual features extracted by models like SyncNet [8, 37]. Although these metrics effectively quantify the quality and synchronization of 2D videos, they are not directly applicable to 3D data, as the latter involve completely different measures and analyses.
@@
Since the release of VOCA [10], the first high-quality open dataset for speech-driven 3D talking heads generation, effectively measuring the quality of the animations has been a challenge. Indeed, this task requires metrics capturing both the temporal aspect (mostly lip synchronization), and the non-Euclidean geometry of meshes in 3D space. To this aim, MeshTalk [43] proposed the Lip-Vertex Error (LVE) that measures the maximum quadratic error between the mouth vertices of the ground truth and predictions. However, LVE focuses on maximum error, which can be misleading: it may be influenced by a single inaccurate vertex, while the overall lip-sync quality remains high. Furthermore, the LVE metric suffers from ambiguity in defining mouth vertices, leading to inconsistent results across different studies due to varying lip vertex masks. This metric was then adopted in most subsequent works. CodeTalker [57], in addition to LVE, proposed the Face Dynamics Deviation (FDD) to measure upper-face movement quality. Despite natural and dynamic face movements are important for realism, the primary focus when generating 3D talking heads should be the accuracy and intelligibility of the speech. FaceDiffuser [47] further introduced two additional metrics: Mean Vertex Error (MVE) and Diversity. MVE measures the maximum of the mean squared error between the ground truth and predictions, while Diversity assesses the variation in generated animations given the same audio and face. While these metrics provide insights into certain aspects of the animation, they still fall short in capturing the true quality of dynamic lip-syncing. To account for the temporal dynamics, Imitator [50] introduced Dynamic Time Warping (DTW) to evaluate the quality of generated faces across all vertices and mouth vertices, yet the ambiguity of the masking factor for the mouth remains. On a similar trend, Nocentini et al[29] proposed the Displacement Angle Discrepancy (DAE), which measures the angular difference between subsequent frames of ground-truth and generated animations. This helped in providing a coarse measure of the consistency of vertices motion. However, it does not tell much about the geometric similarity, as two similar motions might result from very different meshes. On a different perspective, SelfTalk [35] proposed the Lip Readability Percentage (LRP) to assess the intelligibility of spoken words. However, it is an approximate measure based on the distance between mouth vertices. It is basically a thresholded version of the LVE. FaceTalk [2] showcased how other metrics on the rendering of talking heads can be efficiently used such as Lip Sync Error Distance (LSE-D) [37], FID/KID [42], Face Image Quality Assessment [34] and Video Quality Assessment [55]. While providing interesting insights regarding the qualitative performance of the models, they may loose non-linear 3D contexts after projection in flat spaces and are only applcable for textured meshes.

### v3

@@ Abstract
Generating speech-driven 3D talking heads presents numerous challenges; among those is dealing with varying mesh topologies where no point-wise correspondence exists across the meshes the model can animate. While previous literature works assume fixed mesh structures, in this work we present the first framework capable of animating 3D faces in arbitrary topologies, including real scanned data. Our approach leverages heat diffusion to predict features that are robust to the mesh topology. We explore two training settings: a registered one, in which meshes in a training sequences share a fixed topology but any mesh can be animated at test time, and an fully unregistered one, which allows effective training with varying mesh structures. Additionally, we highlight the limitations of current evaluation metrics and propose new metrics for better lip-syncing evaluation. An extensive evaluation shows our approach performs favorably compared to fixed topology techniques, setting a new benchmark by offering a versatile and high-fidelity solution for 3D talking heads where the topology constraint is dropped. The code along with the pre-trained model, is available here.
@@ 1 Introduction
3D talking heads refer to the task of animating a generic 3D face mesh so that it corresponds to a speech, given in the form of an audio signal. This task has received significant attention in recent years, thanks to the wide range of potential applications, across diverse fields, that it can open. This includes computer-generated imagery in movies, video games, virtual reality, and medical simulations. Nevertheless, it also faces unique challenges, primarily arising from the complexity of mapping the semantics of the spoken words contained in the audio to the intricacy of facial motions. A major challenge for 3D talking heads is that of generating convincing and audio-synchronized lip movements that faithfully replicate the spoken sentence. Several other aspects, though, contribute to making a generated animation realistic. For example, head movements, facial expressions or speech styles are all important factors, and numerous approaches were proposed to account for such diverse challenges [nocentini2024emovoca, peng2023emotalk, FaceDiffuser_Stan_MIG2023]. In all cases, the primary objective is that of learning a mapping between the audio input to non-rigid and time-dependent deformations of the face mesh that allow for its animation.
@@
In this regard, all state-of-the-art methods in this field simplify the problem by assuming that the meshes to be animated are aligned to a fixed topology, that is, all face meshes share a consistent number and arrangement of points. On one hand, this constraint allows for eluding most difficulties related to processing unordered point-clouds. With the meshes in vertex-wise correspondence and their structure known and consistent, one can, for instance, use effective graph operators [bouritsas_2019_neural3DMM, chebnet_2016, gong2019spiralnet++] or define specific losses to control localized face areas [Thambiraja_2023_ICCV_imitator]. On the other hand, it puts several constraints on the applicability. For example, existing approaches require training separate models for each specific topology, hindering its generalization power. In addition, if one wants to animate a new mesh, either captured with a real scanner or designed with some graphics software, it is first required to process the mesh to make it consistent with the topology used for training. This also adds a level of complexity in the data collection since newly captured mesh sequences need to be processed in order to align them to a common topology, which unveils new challenges mostly when non-rigid face deformations are involved [ferrari2021sparse, gilani2017dense]. In [nocentini2024scantalk3dtalkingheads], the training process is supervised by the vertex-wise correspondence in the data and the model is robust to generalize to different topologies at inference time. A step further is not assuming vertex-wise correspondence within sequences during the training process (unregistered sequences). In this scenario, training is not supervised by vertex-wise correspondence. This is especially interesting because it allows one to train on minimally processed data, enhancing even more the versatility of the method.
@@
In addition to the aforementioned issues, another critical aspect is related to the evaluation protocols and metrics. Thoroughly assessing the performance of a 3D talking head method is complex, as it requires to simultaneously verify the extent to which: (i) the overall shape of the animated face is preserved. There is in fact no guarantee that mouth motions won’t negatively affect other face regions; (ii) the accuracy of the lip motion and synchronization is maintained, both in terms of shape similarity, perceived realism and motion dynamics. We identified a clear shortcoming in the current literature in this regard. In most works, the primary metrics used for quantitative evaluation and comparison measure the discrepancy between ground-truth and generated animation in terms of per-frame geometric discrepancy of vertices [VOCA2019, Fan_Lin_Saito_Wang_Komura_faceformer_2022, peng2023selftalk, FaceDiffuser_Stan_MIG2023]. These are referred to as Lip-Vertex Error (LVE), Mean Vertex Error (MVE) and Face Dynamics Deviation (FDD). We argue that these metrics do not provide a thorough picture of a method performance, and while these quantitative measures are widely used, they often fall short in providing a comprehensive evaluation. Suffice it to say, the evaluation of the motion dynamics and synchronization is largely ignored, and lower metrics do not necessarily imply a more realistic animation. A common workaround involves performing qualitative user studies, where people are asked to evaluate the quality of the generated samples. While useful to some extent, these studies lack a clear robust protocol, and usually involve a few users and examples, posing doubts on their reliability. Some papers raised similar concerns, and defined additional metrics to fill in such gap. For instance, in [peng2023selftalk] the Lip Readability Percentage (LRP) is introduced. However, it is still based on per-frame lip vertex distances. Other examples are Dynamic Time Warping (DTW) used in Imitator [Thambiraja_2023_ICCV_imitator] for assessing temporal coherence, or the Displacement Angle Discrepancy (DAE) proposed in [landmarks_3D_Nocentini_2023], which provides a measure of the angular difference between subsequent frames. These measures, and others, are a clear evidence that the current benchmarking miss a piece, and that a stable, consistent and comprehensive set of complementary error measures needs to be defined for proper evaluation across methods.
@@
In summary, ScanTalk, illustrated in Figure 1, represents a significant advancement in the field of speech-driven 3D face animation by overcoming the limitations of fixed-topology models and establishing new standards for evaluation and training. Our framework demonstrates broad applicability across different datasets and conditions. Parts of the current work have been introduced in our previous paper [nocentini2024scantalk3dtalkingheads]. The present extension provides the following new contributions:
  • In [nocentini2024scantalk3dtalkingheads], ScanTalk needs to be trained with homogeneous sequences: each frame of the sequence shares the topology of the first frame. Here, we modify the training scheme and define appropriate losses so that the model can be trained with inhomogeneous sequences: each frame of the sequence has an arbitrary topology. By dropping this constraint, ScanTalk can be trained with truly unregistered meshes;
  • While the Chamfer distance is widely used as a loss for unsupervised static 3D reconstruction, we propose to use a dynamic extension of this loss for a complete unsupervised setting to learn speech-driven motion prediction;
@@ 2.1 Animation of Unregistered 3D Face Meshes
Animating a 3D face mesh to reproduce expressions or speech is a research topic that witnessed noticeable advancements only recently, thanks to the progresses achieved in the design of effective deep architectures for processing 3D data [bouritsas_2019_neural3DMM, lu2024fc, otberdout2022sparse, otberdout2023generating]. However, it is well known that deep models are data-hungry, which poses the problem of collecting sufficiently large 4D datasets. Moreover, all state-of-the-art methods based on deep learning require the face meshes to be in a fixed topology.
In the realm of 3D face mesh acquisition for datasets collection, methodologies vary from extracting geometry from images or videos [yi2022_talkshow] or generating synthetic data [principi2023florence], to employing specialized scanners in controlled environments [BIWI_2010, ferrari2023florence, COMA:ECCV18, wuu2022multiface, BU3DFE_2006]. While methods that follow the former paradigm are simpler and cost-effective, and typically result in meshes with known topology, they may not capture complete 3D information with the necessary fidelity. On the other hand, 3D scans, though rich in captured details of the face, often come unregistered. The varying mesh structure and nuisances resulting from the sensors, e.g., holes or noise, further complicate a direct animation. In response, a widely used approach to address this challenge is that of first registering the raw scans onto predefined topologies, typically by means of 3D Morphable Models [BIWI_2010, ferrari2021sparse, FLAME:SiggraphAsia2017, muralikrishnan_2023_BLISS, wuu2022multiface].
While some deep learning approaches have emerged to handle face scans with unknown structure through latent representations with robust encoders [sharp2021diffusion, Wiersma2022DeltaConv] built upon PointNet [Charles_PointNet_2017, qi2017pointnet++] or adaptive convolution strategies [dgcnn], these methods still rely on a registered decoding strategy, which can smooth out important details. Recent advancements closer to our work include DiffusionNet [sharp2021diffusion] combined with neural Jacobian fields [NeuralJacobianField_2022] in Neural Face Rigging (NFR) [Qin_2023_NFR], which facilitates animation transfer between sequences of unregistered face meshes. Our framework stands out by introducing a model that leverages audio data to animate unregistered meshes directly.
@@
Numerous methodologies have emerged to synchronize facial animations with speech. While significant progress has been made in 2D talking heads [EAMM_2022_2Dtalkinghead, Wang2023Emotional_2Dtalkinghead], extending these approaches to 3D faces remains challenging. Nevertheless, generating 3D talking heads from audio has seen significant advancements, initially pioneered by [VOCA2019] and greatly improved through deep learning methods. Recent works leverage large datasets and deep learning strategies [VOCA2019, Fan_Lin_Saito_Wang_Komura_faceformer_2022, facexhubert, landmarks_3D_Nocentini_2023, nocentini2024emovoca, richard2021meshtalk, FaceDiffuser_Stan_MIG2023, Thambiraja_2023_ICCV_imitator, xing2023codetalker]. VOCA [VOCA2019] was among the first to develop a generalizable model for 3D talking heads but lacked expressiveness and convincing lip synchronization. MeshTalk [richard2021meshtalk] improved upon this with a larger registered dataset and a more expressive model. FaceFormer [Fan_Lin_Saito_Wang_Komura_faceformer_2022] utilized a transformer architecture and a pre-trained audio encoder, like Wav2vec2 [wav2vec_2019], with further improvements in FaceXHubert [facexhubert], demonstrating enhanced cross-modality mappings from audio to facial motion. From a different perspective, Nocentini et al. [landmarks_3D_Nocentini_2023] designed a coarse-to-fine approach named S2L-S2D, and showed that lip motion can be effectively modeled by learning to first animate facial landmarks, and then reconstructing the whole face shape. Recent models like CodeTalker [xing2023codetalker] and SelfTalk [peng2023selftalk] introduced new techniques to address over-smoothing and enhance lip motion expressiveness. Integrating emotions [danvevcek2023emotional, peng2023emotalk] and head poses [sun2023diffposetalk, SadTalker_2023] into models further pushes the boundaries, though the lack of registered data remains a significant challenge. In this regard, a recent work attempted to bridge this gap by proposing a data-driven approach to synthetically generate a large-scale dataset of expressive 3D talking heads, named EmoVOCA [nocentini2024emovoca], which combines the VOCAset dataset [VOCA2019] with 4D expressive sequences from the Florence4D dataset [principi2023florence]. However, creating this dataset was made easy as both VOCAset and Florence4D contain meshes in the same topology. More recent approaches leverage denoising diffusion models for generating 3D talking heads with head motions [sun2023diffposetalk, thambiraja2023_3diface, SadTalker_2023], and expressive lip motion generation [chen2025diffusiontalker]. However, these diffusion-based frameworks still suffer from the limitation of fixed topology of the mesh; some of them, such as DiffPoseTalk [sun2023diffposetalk] require a morphable model, thus hindering the generalizability to other face datasets.
@@
In 2D talking head generation, metrics like Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), are commonly employed to evaluate visual fidelity. Meanwhile, lip synchronization can be measured using metrics such as Lip Sync Error Confidence (LSE-C), and Lip Sync Error Distance (LSE-D), which rely on visual features extracted by models like SyncNet [Chung16a_SyncNet, Wav2Lip]. Although these metrics effectively quantify the quality and synchronization of 2D videos, they are not directly applicable to 3D data, as the latter involve completely different measures and analyses.
@@
Since the release of VOCA [VOCA2019], the first high-quality open dataset for speech-driven 3D talking heads generation, effectively measuring the quality of the animations has been a challenge. Indeed, this task requires metrics capturing both the temporal aspect (mostly lip synchronization), and the non-Euclidean geometry of meshes in 3D space. To this aim, MeshTalk [richard2021meshtalk] proposed the Lip-Vertex Error (LVE) that measures the maximum quadratic error between the mouth vertices of the ground truth and predictions. However, LVE focuses on maximum error, which can be misleading: it may be influenced by a single inaccurate vertex, while the overall lip-sync quality remains high. Furthermore, the LVE metric suffers from ambiguity in defining mouth vertices, leading to inconsistent results across different studies due to varying lip vertex masks. This metric was then adopted in most subsequent works. CodeTalker [xing2023codetalker], in addition to LVE, proposed the Face Dynamics Deviation (FDD) to measure upper-face movement quality. Despite natural and dynamic face movements are important for realism, the primary focus when generating 3D talking heads should be the accuracy and intelligibility of the speech. FaceDiffuser [FaceDiffuser_Stan_MIG2023] further introduced two additional metrics: Mean Vertex Error (MVE) and Diversity. MVE measures the maximum of the mean squared error between the ground truth and predictions, while Diversity assesses the variation in generated animations given the same audio and face. While these metrics provide insights into certain aspects of the animation, they still fall short in capturing the true quality of dynamic lip-syncing. To account for the temporal dynamics, Imitator [Thambiraja_2023_ICCV_imitator] introduced Dynamic Time Warping (DTW) to evaluate the quality of generated faces across all vertices and mouth vertices, yet the ambiguity of the masking factor for the mouth remains. On a similar trend, Nocentini et al[landmarks_3D_Nocentini_2023] proposed the Displacement Angle Discrepancy (DAE), which measures the angular difference between subsequent frames of ground-truth and generated animations. This helped in providing a coarse measure of the consistency of vertices motion. However, it does not tell much about the geometric similarity, as two similar motions might result from very different meshes. On a different perspective, SelfTalk [peng2023selftalk] proposed the Lip Readability Percentage (LRP) to assess the intelligibility of spoken words. However, it is an approximate measure based on the distance between mouth vertices. It is basically a thresholded version of the LVE. FaceTalk [aneja2023facetalk] showcased how other metrics on the rendering of talking heads can be efficiently used such as Lip Sync Error Distance (LSE-D) [Wav2Lip], FID/KID [FID_KID_2023], Face Image Quality Assessment [SDD-FIQA2021] and Video Quality Assessment [VQA_wu2023dover]. While providing interesting insights regarding the qualitative performance of the models, they may loose non-linear 3D contexts after projection in flat spaces and are only applcable for textured meshes.